OpenStudy (anonymous):

Find the equation of the tangent line to the curve y=xsqrtx at the point (4,8).

OpenStudy (campbell_st):

well \[x \sqrt{x} = \sqrt{x^3}\] so write it in index form \[y = x^{\frac{3}{2}}\] makes it easy to differentiate

OpenStudy (anonymous):

so the derivative was 3/2 x^1/2

OpenStudy (campbell_st):

yep... thats it... now substitute x = 4 to find the slope at (4, 8)

OpenStudy (anonymous):

so i got for my equation y=3x-4 is that right

OpenStudy (campbell_st):

yep... make sense to me... well done

OpenStudy (anonymous):

thanks